How many aluminum ions are contained in 0.0128 uL of a 2.56 * 10 ^-4 M solution of Al(NO3)3
Answer #1

contained solution of How many aluminium ions are in 0.0128 uL of a 2.56 x Lot M AlCNO3)3 Answer = Number of aluminium ion 0.2.56 X 10 4 M = mole of AICNO3)3 0.0128 x 10-6 L mole of AlCNO3)2 = 2.56* 10 4 x 0.0128 x 16mole mole of AICNO3)3 = 0:032768since we know I mole contains 6.023 X 10 23 atom which is equal to avogadrod number so 1 mole of aluminium ion contains 6.02summary:

Answer = 0.197×10^10 ion

Solution:

First we calculate mole of Al(NO3)^3 using molarity formula.

Mole of Al(NO3)3 = molarity of solution × volume of solution ( in L)

Since volume of solution has unit microlitre. So we first changed the unit of volume in litre .

1uL = 10^ -6 L

After calculating mole of Al(NO3)^3 , we can calculate mole of aluminum ion which is equal to mole of Al(NO3)^3.

After calculating mole of Aluminium ion , we multiply mole of Aluminium ion by 6.023 ×10^23. So we get number of Aluminium ion.